Solve the equation by extracting square roots:
(x +5)^2 = (x + 4)^2
My answer is x = -9/2 and I am not sure it is correct. I squared both sides and then used x-5 on each side of the equation and -x-4.
left side was 2x = -9 then I divided 2 into each side leaving x = -9/2. Is this the correct way to complete this equation? and is this the correct answer/

Answer by checkley75(3666)

(x+5)^2=(x+4)^2 squaring both sides
x^2+10x+25=x^2+8x+16 now combine like terms
x^2+x^2+10x-8x=16-25 reduce
2x=-9
x=-9/2
x=-4.5